\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}
\providecommand{\MR}{\relax\ifhmode\unskip\space\fi MR }
% \MRhref is called by the amsart/book/proc definition of \MR.
\providecommand{\MRhref}[2]{%
  \href{http://www.ams.org/mathscinet-getitem?mr=#1}{#2}
}
\providecommand{\href}[2]{#2}
\begin{thebibliography}{DDT94}

\bibitem[Aga]{congnum}
A.~Agashe, \emph{The modular number, the congruence number, and multiplicity
  one}, preprint, available at \hfill \\ {\sf http://www.math.fsu.edu/\~{
  }agashe/congnum.html}.

\bibitem[Aga13]{conjs}
Amod Agashe, \emph{Conjectures concerning the orders of the torsion subgroup,
  the arithmetic component groups, and the cuspidal subgroup}, Exp. Math.
  \textbf{22} (2013), no.~4, 363--366. \MR{3171097}

\bibitem[Aga18]{tor}
\bysame, \emph{Rational torsion in elliptic curves and the cuspidal subgroup},
  J. Th\'{e}or. Nombres Bordeaux \textbf{30} (2018), no.~1, 81--91.
  \MR{3809710}

\bibitem[BD14]{bd}
Nicolas Billerey and Luis~V. Dieulefait, \emph{Explicit large image theorems
  for modular forms}, J. Lond. Math. Soc. (2) \textbf{89} (2014), no.~2,
  499--523. \MR{3188630}

\bibitem[Cre97]{cremona:algs}
J.\thinspace{}E. Cremona, \emph{Algorithms for modular elliptic curves}, second
  ed., Cambridge University Press, Cambridge, 1997.

\bibitem[DDT94]{ddt}
H.~Darmon, F.~Diamond, and R.~Taylor, \emph{Fermat's last theorem}, Current
  developments in mathematics, 1995 (Cambridge, MA), Internat. Press,
  Cambridge, MA, 1994, pp.~1--154.

\bibitem[DR73]{deligne-rapoport}
P.~Deligne and M.~Rapoport, \emph{Les sch\'emas de modules de courbes
  elliptiques}, Modular functions of one variable, II (Proc. Internat. Summer
  School, Univ. Antwerp, Antwerp, 1972) (Berlin), Springer, 1973, pp.~143--316.
  Lecture Notes in Math., Vol. 349.

\bibitem[Edi92]{edixhoven:weight}
B.~Edixhoven, \emph{The weight in \protect{S}erre's conjectures on modular
  forms}, Invent. Math. \textbf{109} (1992), no.~3, 563--594.

\bibitem[Kat73]{katz:antwerp350}
N.\thinspace{}M. Katz, \emph{$p$-adic properties of modular schemes and modular
  forms}, Modular functions of one variable, III (Proc. Internat. Summer
  School, Univ. Antwerp, Antwerp, 1972) (Berlin), Springer, 1973, pp.~69--190.
  Lecture Notes in Mathematics, Vol. 350.

\bibitem[Maz77]{mazur:eisenstein}
B.~Mazur, \emph{Modular curves and the \protect{Eisenstein} ideal}, Inst.
  Hautes \'Etudes Sci. Publ. Math. (1977), no.~47, 33--186 (1978).

\bibitem[Maz78]{mazur:rational}
\bysame, \emph{Rational isogenies of prime degree (with an appendix by {D}.
  {G}oldfeld)}, Invent. Math. \textbf{44} (1978), no.~2, 129--162.

\bibitem[Ogg73]{ogg}
A.~P. Ogg, \emph{Rational points on certain elliptic modular curves}, Analytic
  number theory ({P}roc. {S}ympos. {P}ure {M}ath., {V}ol. {XXIV}, {S}t. {L}ouis
  {U}niv., {S}t. {L}ouis, {M}o., 1972), Proc. Sympos. Pure Math., vol. Vol.
  XXIV, Amer. Math. Soc., Providence, RI, 1973, pp.~221--231. \MR{337974}

\bibitem[Oht14]{ohta:eisenstein}
Masami Ohta, \emph{Eisenstein ideals and the rational torsion subgroups of
  modular {J}acobian varieties {II}}, Tokyo J. Math. \textbf{37} (2014), no.~2,
  273--318. \MR{3304683}

\bibitem[Rib90]{ribet:modreps}
K.\thinspace{}A. Ribet, \emph{On modular representations of \protect{${\rm
  {G}al}(\overline{\bf{Q}}/{\bf {Q}})$} arising from modular forms}, Invent.
  Math. \textbf{100} (1990), no.~2, 431--476.

\bibitem[Rib08]{ribet:eisenstein}
K.~A. Ribet, \emph{Eisenstein primes for~{$J_0(pq)$}}, unpublished (2008).

\bibitem[Ser72]{serre:propgal}
J-P. Serre, \emph{Propri\'et\'es galoisiennes des points d'ordre fini des
  courbes elliptiques}, Invent. Math. \textbf{15} (1972), no.~4, 259--331.

\bibitem[Shi94]{shimura:intro}
G.~Shimura, \emph{Introduction to the arithmetic theory of automorphic
  functions}, Princeton University Press, Princeton, NJ, 1994, Reprint of the
  1971 original, Kan Memorial Lectures, 1.

\bibitem[Ste]{stein:cuspidal}
W.\thinspace{}A. Stein, \emph{The cuspidal subgroup of {$J_0(N)$},\hfill\\ {\tt
  https://wstein.org/tables/cuspgroup/index.html}}.

\bibitem[Tak97]{takagi:cuspidal}
Toshikazu Takagi, \emph{The cuspidal class number formula for the modular
  curves {$X_0(M)$} with {$M$} square-free}, J. Algebra \textbf{193} (1997),
  no.~1, 180--213. \MR{1456573}

\bibitem[Tan97]{tang:congruences}
Shu-Leung Tang, \emph{Congruences between modular forms, cyclic isogenies of
  modular elliptic curves and integrality of {$p$}-adic {$L$}-functions},
  Trans. Amer. Math. Soc. \textbf{349} (1997), no.~2, 837--856.

\bibitem[Vat05]{vatsal:mult}
V.~Vatsal, \emph{Multiplicative subgroups of {$J\sb 0(N)$} and applications to
  elliptic curves}, J. Inst. Math. Jussieu \textbf{4} (2005), no.~2, 281--316.

\bibitem[Yoo16]{yoo2}
Hwajong Yoo, \emph{On {E}isenstein ideals and the cuspidal group of
  {$J_0(N)$}}, Israel J. Math. \textbf{214} (2016), no.~1, 359--377.
  \MR{3540618}

\bibitem[Yoo19]{yoo:non-optimal}
\bysame, \emph{Non-optimal levels of a reducible {$\rm{mod}\,\ell$} modular
  representation}, Trans. Amer. Math. Soc. \textbf{371} (2019), no.~6,
  3805--3830. \MR{3917209}

\bibitem[Yoo23]{yoo:rational}
\bysame, \emph{The rational cuspidal divisor class group of {$X_0(N)$}}, J.
  Number Theory \textbf{242} (2023), 278--401. \MR{4490453}

\end{thebibliography}

\documentclass[11pt]{article}

\makeindex
\usepackage[all]{xy}
\usepackage{graphicx}

\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{amsthm}

\newcommand{\isom}{\cong}

\newcommand{\Z}{{\bf{Z}}}

\newcommand{\Q}{{\bf{Q}}}

\newcommand{\R}{{\bf{R}}}
\newcommand{\C}{{\bf{C}}}

\newcommand{\T}{{\bf{T}}}

\newcommand{\m}{{\mathfrak{m}}}

\newcommand{\Mid}{|} 

\newcommand{\miD}{|}

\newcommand{\divs}{\!\mid\!}
\newcommand{\ndiv}{\!\nmid\!}
\newcommand{\tensor}{\otimes}

\newcommand{\ra}{{\rightarrow}}

 1
\DeclareFontEncoding{OT2}{}{} 
  \newcommand{\textcyr}[1]{%
    {\fontencoding{OT2}\fontfamily{wncyr}\fontseries{m}\fontshape{n}%
     \selectfont #1}}
\newcommand{\Sha}{{\mbox{\textcyr{Sh}}}} 

\newcommand{\pp}{{\mathfrak{p}}}

\newcommand{\OO}{{\mathcal{O}}}
\newcommand{\Of}{{\OO_{\scriptscriptstyle{f}}}}

\usepackage{bm}
\bmdefine\bmu{\mu}

\newcommand{\comment}[1]{}

\DeclareMathOperator{\tor}{tor}

\newtheorem{lem}{Lemma}[section]
\newtheorem{cor}[lem]{Corollary}
\newtheorem{prop}[lem]{Proposition}
\newtheorem{conj}[lem]{Conjecture}
\newtheorem{thm}[lem]{Theorem}

\theoremstyle{definition}

\newtheorem{rmk}[lem]{Remark}

\newcommand{\thetitle}
{Reducibility and rational torsion in modular abelian varieties}

\begin{document}
\parindent=2em

\title{\thetitle}
\author{Amod Agashe and Matthew Winters}

\date{}
\maketitle

\begin{abstract}
Let $N$ be a square-free positive integer and let
$f$ be a newform of weight~$2$ on~$\Gamma_0(N)$. 
Let $A$ denote the abelian subvariety of~$J_0(N)$ associated to~$f$ and
let ${\m}$ be a maximal ideal of the Hecke algebra that contains 
${\rm Ann}_{\T} f$ and has residue characteristic~$r$ such that
$r \ndiv 6N$. 
We show that if either $A[{\m}]$ or
the canonical representation~$\rho_{{\m}}$ over~$\T/\m$ associated to~$\m$
is reducible, then $r$ divides the order of the cuspidal subgroup of~$J_0(N)$
and $A[{\m}]$ has a nontrivial rational point. We mention some 
applications
of this result, including an application 
to the second part of the Birch and Swinnerton-Dyer 
conjecture for~$A$.
\end{abstract}

\section{Introduction} \label{intro}
Let $N$ be a positive integer. Let
$X_0(N)$ denote the modular curve over~$\Q$ associated to~$\Gamma_0(N)$,
and let $J_0(N)$ denote its Jacobian, which is an abelian variety
over~$\Q$.
Let $f$ be a newform of weight~$2$ on~$\Gamma_0(N)$. 
Let $\T$ denote the subring of endomorphisms of~$J_0(N)$
generated by the Hecke operators (usually denoted~$T_\ell$
for $\ell \ndiv N$ and $U_p$ for $p\divs N$). 
Let $I_f = {\rm Ann}_{\T} f$ and let 
$A$ denote the abelian subvariety of~$J_0(N)$ associated to~$f$
(see~\cite[Theorem~7.14]{shimura:intro}).

If $f$ has integer Fourier coefficients, then $A$ is just an elliptic curve,
and it is said to be {\em optimal}; 
suppose
that is the case for the moment. 
Let $r$ be a prime number.
If $A$ has a nontrivial rational point of order~$r$, then clearly
$A[r]$ is reducible
as a representation of~${\rm Gal}(\overline{\Q}/\Q)$. The converse
need not be true.
For example,
for $A = 99d1$ (the notation is as in~\cite{cremona:algs}), $A[5]$ is reducible, but $A$ has no
nontrivial rational $5$-torsion. 
However, the first author had conjectured:

\begin{conj}\cite[Conjecture~2.5]{conjs} \label{conj1}
Suppose $N$ is square-free, i.e., $A$ is semistable, 
and $r$ is an odd prime. 
If $A[r]$ is
reducible, then $A$ has a nontrivial rational~$r$-torsion point 
(recall that we are assuming that $A$ is optimal).
\end{conj}

Note that if $A[2]$ is reducible, then being two dimensional, 
it has a one dimensional invariant subspace, which has to be of the form 
$\{O,P\}$ for some point~$P$; 
the Galois group has to preserve~$O$ and any element $\sigma$ of it cannot move~$P$ to~$O$, 
since its inverse would move~$O$ to~$P$, and thus $\sigma$ has to preserve~$P$, 
making~$P$ a rational point. So the hypothesis that~$r$ 
is odd could have been dropped in the conjecture above, but is 
relevant for a suitable generalization of the conjecture to more general 
abelian varieties associated to newforms (where the argument above does not work).

The conjecture above follows 
from~\cite{vatsal:mult}, though it is not stated as a result there,
since loc. cit. is primarily concerned with other topics.
For the benefit of the literature, we state it as a theorem:

\begin{thm}[Vatsal] \label{vatsal}
If $A$ is an optimal 
semistable elliptic curve
and $r$ is a prime such that 
$A[r]$ is reducible, then $A$ has a nontrivial rational $r$-torsion point.
\end{thm}
\begin{proof}
If $r$ is odd, then by Lemma~5.2 of~\cite{vatsal:mult}, $A$ has ordinary reduction at~$r$; using
this, Proposition~5.3(ii) of~\cite{vatsal:mult} tells us that 
$A$ has a nontrivial rational~$r$-torsion point.
If $r=2$, then as argued above, the conclusion holds (even without the 
optimality or semistability
hypotheses).
\end{proof}

We remark that the optimality hypothesis cannot be removed in the theorem above: 
the elliptic curve $A = 26b2$
has no nontrivial rational $7$-torsion, while $A[7]$ is reducible; here $A$ is not optimal.

In~\cite{conjs}, the author had  remarked that
it was hoped that Conjecture~\ref{conj1} above
``might be  appropriately generalized
to quotients or subvarieties of arbitrary dimension of~$J_0(N)$ associated to newforms'';
at that time, lacking numerical evidence beyond the elliptic curve case, the author
did not make such a generalization.
Let us revert to the case where $f$ need not have integer Fourier coefficients,
so that $A$ need not be an elliptic curve.
It is not immediately clear to the authors of this article if the techniques of~\cite{vatsal:mult}
generalize to this case, since it is a rather lenghty article, whose main goal
is something else. It would be interesting to try to generalize everything in 
that article beyond the elliptic curve case, which would be a bigger project;
in any case, it is not clear a priori what would generalize.

Let $C$ denote the cuspidal subgroup of~$J_0(N)$. If
$N$ is square-free, then all the points of~$C$ are rational
(e.g., by~\cite[Prop.~2]{ogg}).
Recall that to a maximal ideal~$\m$ of the Hecke algebra, there is associated
a canonical Galois representation~$\rho_{\m}: {\rm Gal}(\overline{\Q}/\Q)
\ra {\rm GL}_2(\T/\m)$
(see~\cite[Prop.~5.1]{ribet:modreps}). 
In this article, we prove:
\begin{thm}\label{thm:main}
Suppose $N$ is square-free and ${\m}$ is a maximal ideal of~$\T$ 
that contains~$I_f$ and has 
residue characteristic~$r$ such that $r \ndiv 6N$.
If either $\rho_{\m}$ or~$A[\m]$ is reducible as a representation over~$\T/\m$, then
$A[\m] \cap C$ is nontrivial, and so\\
(a) $r$ divides $|C|$, \\
(b) $A[\m]$ has a nontrivial rational point, and \\
(c) $A$ has a nontrivial rational $r$-torsion point. \\
In particular, if 
$A$ does not have any nontrivial rational $r$-torsion point
or $r$ does not divide $|C|$, then 
$\rho_{\m}$ is irreducible. 
\end{thm}

The proof of the theorem above is given in Section~\ref{sec:lemmas}.
The idea of the proof is very similar 
to that of the main theorem of~\cite{tor}: we use the hypotheses to 
show that the newform~$f$ 
is congruent to an Eisenstein series~$E$
modulo a prime ideal~$\pp$ in the number field generated by the 
Fourier coefficients of~$f$ that lies over~$r$ 
(the tricky part is to prove the congruences for Fourier coefficients
of indices that are not coprime to~$N$). 
As part of the proof of this congruence,
we show that under the hypotheses of the theorem,
for at least one prime~$p$ that divides~$N$,
the sign of the Atkin-Lehner involution at~$p$ acting on~$f$ is~$-1$, which 
is an interesting result on its own
(see Proposition~\ref{prop:mazur2}; some versions of this result have
already appeared in the literature: see Remark~\ref{rmk2}).
Given the congruence between~$f$ and~$E$, and the fact that
$f$ is ordinary at~$\pp$ (which we show), a result of 
Tang~\cite[Thm~0.4]{tang:congruences}
tells us that 
$A[{\m}]$ has nontrivial
intersection with 
a subgroup of the cuspidal group~$C$ of~$J_0(N)$, giving us the theorem. 

The hypothesis that $\rho_{\m}$ is irreducible is often used in arithmetic geometry
(for example in multiplicity one results, as in~\cite[Prop.~1.3]{congnum}). 
One significance of the theorem above is that
the condition that 
$A$ does not have any nontrivial rational $r$-torsion point
may  be easier to check
theoretically or in numerical examples, 
than whether $\rho_{\m}$ is irreducible. 
Another significance of the theorem stems from the fact that the cuspidal subgroup
can be computed (see, e.g.~\cite{stein:cuspidal}), and the 
structure of~$C$ when $N$ is square-free
is given in~\cite{takagi:cuspidal} (see also~\cite{yoo:rational}; the formulas
are too complicated to mention here). As an 
example, the order of the cuspidal subgroup for $N=14$ is~$6$, so for any
maximal ideal~$\m$ of~$\T$ (of level~$14$) of residue characteristic bigger than~$7$, 
$\rho_{\m}$ is irreducible. 

In the next section, we mention 
some more applications of Theorems~\ref{vatsal} and~\ref{thm:main}, 
including one relevant to the second part of the Birch and Swinnerton-Dyer conjecture.
In Section~\ref{sec:lemmas}, we prove some results regarding the 
Fourier coefficients of~$f$ that are needed to show the congruence
alluded to above and use these
results to prove Theorem~\ref{thm:main}; the reader who is interested
in these results and the proof may go directly to that section.\\

\noindent {\em Acknowledgement:}
We are grateful to E.~Ghate for some discussions regarding this article.

\section{Applications} \label{sec:app}
We continue to use the notation introduced in the previous section.

By a theorem of Mazur~(\cite[Theorem~III.5.1]{mazur:eisenstein}), if $r$ a prime bigger than~$7$,
then $r$ does not divide the order of the torsion subgroup of any elliptic curve over~$\Q$. 
So by Theorem~\ref{vatsal}, we get:
\begin{cor}
Suppose $A$ is an optimal semistable elliptic curve 
and $r$ a prime such that $r> 7$. 
Then $A[r]$ is irreducible. 
\end{cor}
This result is already known by~\cite[Theorem~4]{mazur:rational}; however,
the proof is different. In this context, we should also point out that
if one drops the hypothesis that $N$ is square-free, then 
there is a finite list of primes~$r$ such that an elliptic 
curve has a rational isogeny of degree~$r$: 
see~\cite[Theorem~1]{mazur:rational}.

\comment{
Reverting back to the general case where $A$ need not be an elliptic curve,
there is no known analog of Mazur's theorem on the possible orders of torsion subgroups
mentioned above. However, Ohta has generalized some of Mazur's techniques
in~\cite{mazur:eisenstein} 
to show that if $N$ is square-free and $r > 3$, then the $r$-power torsion
subgroup of~$J_0(N)(\Q)$ is contained in the cuspidal subgroup~$C$
(see the main theorem of~\cite{ohta:eisenstein}). So we get

\begin{prop} \label{irred}
Suppose $N$ is square-free and $\m$ is a maximal ideal of~$\T/I_f$ of
residue characteristic~$r$ such that $r \ndiv 6N$ and $r \ndiv |C|$.
Let $\m$ denote the inverse image of~${\m'}$ under the map $\T \ra \T/I_f$.
Then $\rho_{\m}$ and~$A[{\m'}]$ are both irreducible as representations over~$\T/\m$. 
\end{prop}

The importance of the proposition above stems from the fact that the cuspidal subgroup
can be computed (see, e.g.~\cite{stein:cuspidal}) and the 
structure of~$C$ when $N$ is square-free
is discussed in~\cite{takagi:cuspidal} (see also~\cite{yoo:rational}). We remark
that the proposition above follows directly from the proof of Theorem~\ref{thm:main}
(without having to use~\cite{ohta:eisenstein}); however, we wanted to
emphasize that any bound on the order of the rational torsion subgroup of~$A$
will have a consequence for the irreducibilty of~$\rho_{\m}$ (when the hypotheses
of the theorem hold).

}

As mentioned in~\cite[p.~307]{serre:propgal}, 
if $A$ is a semistable elliptic curve (it need not be optimal) and $r$ is a prime
such that $A[r]$ is reducible 
as a representation of~${\rm Gal}(\overline{\Q}/\Q)$, then 
this representation can be put in matrix form as
$$ 
\begin{bmatrix}
\chi' & * \\
0 & \chi'' 
\end{bmatrix},
$$
where $\chi'$ and $\chi''$ are characters of~${\rm Gal}(\overline{\Q}/\Q)$
such that one of them is the identity character and the other is the \mbox{$\bmod\ r$} cyclotomic character. If $\chi'$ is the identity
character, then $A$ has a nontrivial rational point of order~$r$; however, 
one does not know if $\chi'$ is the identity
character (as per loc. cit.). 
Theorem~\ref{vatsal} above implies that under the additional hypothesis that 
$A$ is optimal, one may take $\chi'$ to be the identity character.

In the rest of this section, we discuss an application of 
Theorems~\ref{vatsal} and~\ref{thm:main}
to the second part of 
the Birch and Swinnerton-Dyer (BSD)
conjecture for~$A$. 
Let $L_A(s)$ denote
the $L$-function of~$A$. 
Let $K_A$ denote the coefficient of the leading term of the Taylor
series expansion of~$L_A(s)$ at $s=1$, and let $R_A$ denote the 
regulator of~$A$.
Let $\Omega_A$ denote the volume of~$A({\R})$ calculated using
a generator
of the group of invariant differentials on the N\'{e}ron 
model of~$A$. 
Let $\Sha_{A}$ denote the Shafarevich-Tate group of~$A$, which
we assume is finite. 
If $p$ is a prime that divides~$N$, then
let $c_p(A)$ denote the order of the arithmetic component group
of~$A$ at~$p$ (also called the Tamagawa number of~$A$ at~$p$).
Then the {\it second part of the BSD~conjecture}
asserts the formula: 
\begin{eqnarray} \label{bsd}
\frac{K_A}{\Omega_A \cdot R_A}
\stackrel{?}{=} 
\frac{\Mid \Sha_{A} \miD \cdot \prod_p c_p(A)}{ \Mid A({\Q})_{\tor} \miD 
\cdot \Mid A^\vee({\Q})_{\tor} \miD }\ .
\end{eqnarray}
Based on numerical evidence, the first author has conjectured:
\begin{conj}\cite[Conjecture~2.4]{conjs} \label{conj2} 
Suppose that $A$ is an elliptic curve (recall that it is optimal by assumption).
If an odd prime~$\ell$ divides $c_p(A)$ for some prime~$p$ that divides~$N$, 
then
either $\ell$ divides $| A(\Q)_{\tor}|$
or the newform~$f$ 
is congruent to a newform of level dividing~$N/p$
(for all Fourier coefficients whose indices are coprime to~$N \ell$)
modulo a prime ideal over~$\ell$ in a number field containing
the Fourier coefficients of both newforms.
\end{conj}
This conjecture indicates some (conjectural) cancellation between the numerator
and denomintor of the right side of the BSD formula~(\ref{bsd}) above
(for more on such cancellations, see~\cite{conjs}).
Towards the conjecture above, we have the following mild generalization 
of~\cite[Proposition~2.3]{conjs}:
\begin{prop} \label{mild}
Let $\ell$ be an odd prime such that 
either $\ell \nmid N$ or for all primes $r$ that divide~$N$,
$\ell \nmid (r-1)$. 
If $\ell$ divides the order of the geometric component
group of~$A$ at~$p$ for some prime~$p || N$,
then either for some maximal ideal~$\m$ of~$\T$ of residue
characteristic~$\ell$ and containing~$I_f$, the representation
$\rho_{\m}$ is reducible, or the newform~$f$ 
is congruent to a newform of level dividing~$N/p$
(for all Fourier coefficients whose indices are coprime to~$N \ell$)
modulo a prime ideal over~$\ell$ in a number field containing
the Fourier coefficients of both newforms.
\end{prop}
\begin{proof}
The proof of Proposition~2.3 in~\cite{conjs} works with the following changes:
replace $E$ by~$A$ and $E[\ell]$ by~$\rho_{\m}$.
\end{proof}
Note that if $p$ is a prime that divides~$N$, and if a prime~$\ell$
divides~$c_p(A)$, then $\ell$ also divides
the order of the geometric component
group of~$A$ at~$p$.

In view of Theorems~\ref{vatsal} and~\ref{thm:main}, from the proposition above, we get:
\begin{prop} \label{prop0}
Let $\ell$ be a prime. Suppose $N$ is square-free and one of the following is true:\\
(a) $\ell \nmid 6N$. \\
(b)  $A$ is an ellitpic curve and 
either $\ell \nmid N$ or for all primes $r$ that divide~$N$, $\ell \nmid (r-1)$. \\
If $\ell$ divides the order of the geometric component
group of~$A$ at~$p$ for some prime~$p \divs N$,
then either $\ell$ divides $| A(\Q)_{\tor}|$
or the newform~$f$ 
is congruent to a newform of level dividing~$N/p$
(for all Fourier coefficients whose indices are coprime to~$N \ell$)
modulo a prime ideal over~$\ell$ in a number field containing
the Fourier coefficients of both newforms.
\end{prop}

The proposition above is a better result towards Conjecture~\ref{conj2}
than Proposition~\ref{mild}, at least 
when $\ell \ndiv 3N$.
Since, as mentioned earlier, if $r$ a prime bigger than~$7$ and $A$ is an elliptic curve,
then $r$ cannot not divide~$| A(\Q)_{\tor}|$, we get from the proposition above:
\begin{cor} Suppose $N$ is square-free and $A$ is an elliptic curve (recall 
that it is optimal by assumption).
Let $\ell$ be a prime such that $\ell > 7$. 
If $\ell$ divides the order of the geometric component
group of~$A$ at~$p$ for some prime~$p \divs N$, then
the newform~$f$ 
is congruent to a newform of level dividing~$N/p$
(for all Fourier coefficients whose indices are coprime to~$N \ell$)
modulo a prime ideal over~$\ell$ in a number field containing
the Fourier coefficients of both newforms.
\end{cor}

Finally, from Theorem~\ref{thm:main} and Proposition~\ref{prop0},  we have:
\begin{cor}
Suppose $N$ is square-free.
Let $\ell$ be a prime such that $\ell \nmid 6N$ and
$\ell$
does not divide~$|C|$.
If $\ell$ divides the order of the geometric component
group of~$A$ at~$p$ for some prime~$p \divs N$,
then 
the newform~$f$ 
is congruent to a newform of level dividing~$N/p$
(for all Fourier coefficients whose indices are coprime to~$N \ell$)
modulo a prime ideal over~$\ell$ in a number field containing
the Fourier coefficients of both newforms.
\end{cor}

\section{Some results on Fourier coefficients and
proof of the main theorem} \label{sec:lemmas}

We continue to use the notation introduced in Section~\ref{intro}.
Let $a_n = a_n(f)$ denote the $n$-th Fourier coefficient of~$f$.
Let $K_f$ denote the number field generated by the Fourier coefficients of~$f$.
We have a ring homomorphism $\T \ra K_f$ that takes $T_n$ to $a_n(f)$. Its
image is an order in~$K_f$, hence it is contained in the 
ring of integers of~$K_f$, which we denote by~$\Of$. 
We thus have an injection $\phi: \T_f := \T/I_f \ra \Of$. 
Recall that  $\m$ is a maximal ideal of~$\T$ that contains~$I_f$ and 
$r$ is the characteristic of~$\T/\m$. Let ${\m'}$ denote
the maximal ideal of~$\T_f$ corresponding to~$\m$.
Since $\Of$ is an integral extension of~$\T_f$, there
is a prime ideal~$\pp$ of~$\Of$ that lies over~${\m'}$, i.e.,
$\T_f \cap \phi^{-1}(\pp) = {\m'}$. In particular, $\phi({\m'}) \subseteq \pp$.
Since $\Of$ is Dedekind, $\pp$ is in fact a maximal ideal, but that will
not be needed in the sequel.
Recall that $\rho_{\m}$ 
denotes the canonical representation associated to~$\m$ (see~\cite[Prop.~5.1]{ribet:modreps}). 
Suppose that $\rho_{\m}$ is reducible.
We do not yet assume the hypotheses in our main theorem that 
$N$ is square-free or $r \ndiv 6N$. 
If $p$ is a prime that divides~$N$, then
let $w_p$ denote the sign of the Atkin-Lehner involution~$W_p$
acting on~$f$.

The following lemma is perhaps well known.
\begin{lem} \label{lem:mazur0}
For all primes $\ell \ndiv N$, we have
$a_\ell(f) \equiv 1 + \ell \bmod \pp$ and 
for all primes $p \divs N$, we have $a_p(f) = - w_p$.
\end{lem}
\begin{proof}
Suppose $\ell \ndiv N$. 
Since $\rho_{\m}$ is reducible, it follows from~\cite[p.~362]{yoo2} that 
$T_\ell - \ell -1 \in \m$. Thus the image of~$T_\ell - \ell -1$ under
the composite $\T/{\m} \ra \T_f/{\m'} \ra \Of/\pp$ is zero; but
this image is the coset represented by 
$a_\ell(f) - 1 - \ell$; hence $a_\ell(f) - 1 - \ell \in \pp$, which
proves the first assertion.

If $p \divs N$, then $a_p(f) = - w_p$ because
$U_p = - W_p$ on the new subspace of~$S_2(\Gamma_0(N),\C)$.
This finishes the proof of the lemma.
\end{proof}

The following proposition is well known, but we give details for the 
convenience of the reader.
\begin{prop}\label{prop:eis}
Suppose $N$ is square-free. 
For every prime~$p$ that divides~$N$,
suppose we are given
an integer $\delta_p \in \{1,p\}$ such that  
$\delta_p = 1$ for at least one~$p$.
Then there is an Eisenstein series~$E$ of weight~$2$ on~$\Gamma_0(N)$ 
which is an eigenfunction for all the Hecke operators such that
for all primes~$\ell \ndiv N$, we have 
$a_\ell(E) = \ell + 1$, and for all primes~$p \divs N$,
we have $a_p(E) = \delta_p$. 
Moreover, $E$ has integer Fourier coefficients, except perhaps~$a_0(E)$,
which is a rational number 
whose denominator (in reduced form) divides~$24$,
and:\\
(a) if for at least one prime~$p$ that divides~$N$ we have
$\delta_p = p$, then $a_0(E) =0$, and\\
(b) if $N$ is prime, then $a_0(E) = (1-N)/24$. 
\end{prop}
\begin{proof}
Everything except the last sentence is~\cite[Proposition~2.1]{tor}.
We now prove the last sentence. Note that $e(q)$ in 
 the first sentence in the proof of Proposition~2.1 in~\cite{tor}
has integer Fourier coefficients
except for the zeroth, which is $1/24$, and since $E$ is obtained
by ``raising levels'' from~$e(q)$, it follows that
$E$ has integer Fourier coefficients, except perhaps~$a_0(E)$,
which is a rational number 
whose denominator (in reduced form) divides~$24$.

We now prove part~(a). Suppose that $\delta_p = p$ for a prime~$p$ that divides~$N$. 
Note that in the proof of Proposition~2.1 in loc. cit.,
in the second last sentence of the second last paragraph, it
says that the $s$-th Fourier coefficient may be chosen to be~$1$ or~$s$;
this is done by using the construction in the second paragraph of the proof in loc. cit.,
with $g_r$ or $\tilde g_r$ respectively. We take $s$ in 
the second last paragraph of the proof of
Proposition~2.1 in loc. cit. to be our~$p$
for which $\delta_p = p$, and 
choose the $s$-th Fourier coefficient to be~$s$; 
so we are using the $\tilde g_r$ construction.
But $\tilde g_r(q) = g(q) - g(q^r)$, and so 
$a_0(\tilde g_r(q)) = a_0(g(q)) - a_0(g(q^r)) 
= a_0(g(q)) - a_0(g(q)) =0$, and so the zeroth Fourier coefficient 
at the end of the second last paragraph of the proof of
Proposition~2.1 in loc. cit. is~$0$.
Further level raising in the last paragraph of that proof does not change the 
zeroth Fourier coefficient, which proves part~(a) of the proposition.

Finally, we prove part~(b). Suppose that $N$ is prime.
Then by hypothesis, $\delta_N = 1$, and so 
we use
the construction~$g_r(q) = g(q) - r \cdot g(q^r)$, with $r=N$ and $g(q)=e(q)$. So
$a_0(E) = a_0(e_N) = (1-N)a_0(e) = (1-N)/24$,
which proves 
part~(b) of the proposition.
\end{proof}

Keeping in mind the strategy of the proof of 
our main theorem (Theorem~\ref{thm:main})
mentioned 
in the introduction, 
we see from the lemma and proposition above 
that coming up with an Eisenstein series~$E$ such that
$a_\ell(f) \equiv a_\ell(E) \bmod r$ for all primes $\ell \ndiv N$
is rather easy. Proving the congruence for all primes~$p$ that divide~$N$
for a suitable Eisenstein  series
is the tricky part, for which we need the results below.

\begin{lem}[Yoo] \label{lem:dum}
Suppose $r \geq 3$.
If $p$ is a prime such that $p \divs N$, $p \neq r$, and $w_p = 1$, then
$p \equiv -1 \bmod \pp$.
\end{lem}

\begin{proof} 
The argument is essentially given in the proof 
of~\cite[Lemma 2.1]{yoo2}; we repeat it here for the convenience of the reader.
Since $\rho_{\m}$ is reducible, 
$\rho_{\m} \isom {\bf 1} \oplus \chi_r$, where ${\bf 1}$ is the
trivial character and $\chi_r$ is the mod~$r$ cyclotomic character.
On the other hand, the semisimplification of the restriction of~$\rho_{\m}$
to~${\rm Gal}(\overline{\Q}_p/\Q_p)$ is isomorphic
to~$\epsilon \oplus \epsilon \chi_r$, where $\epsilon$ is the
unramified quadratic character with $\epsilon({\rm Frob})_p = a_p = -1$
because ${\m}$ is~$p$-new (cf.~\cite[Theorem~3.1(e)]{ddt}). 
From this, we get $p \equiv -1 \bmod r$, and since $r \in \pp$, we get our result. 
\end{proof}

\begin{lem}\label{lem:r}
Suppose $r \geq 3$ and $r \divs N$. Then $w_r = -1$. 
\end{lem}
\begin{proof}
As mentioned on p.~363 in~\cite{yoo2}, we have
$U_r \equiv 1 \bmod {\m}$ by~\cite[Lemma 1.1]{ribet:eisenstein}; it is
mentioned on p.~362 in~\cite{yoo2} that the quoted lemma also 
follows from the result by Deligne given in~\cite[Theorem~2.5]{edixhoven:weight}.
So if $w_r = 1$, then $a_r = -1$, and so
$-1 \equiv 1 \bmod {\m}$, i.e. $2 \in {\m}$, which
is not possible since $r$ is odd; thus $w_r = -1$. 
\end{proof}


\begin{prop} \label{prop:mazur2}
Recall that we are assuming that $\rho_{\m}$ is reducible.
Suppose that $N$ is square-free and
$r > 3$ or that $r=3$ and $\Of/\pp \isom \Z/3\Z$ (the latter happens if $f$ has
rational Fourier coefficients, i.e., $A$ is an elliptic curve). 
Then there is a prime~$p$ that divides~$N$ such that $w_p = -1$, i.e., $a_p(f)=1$.
\end{prop}

\begin{rmk} \label{rmk2}
Special cases of the proposition above have already appeared in the literature.
The result is stated for $r >3$ as Theorem~1.2(1) in~\cite{yoo:non-optimal}, and 
attributed to Ribet, but for a proof, the author refers to Theorem~2.6(ii)(b)
of~\cite{bd};  however in loc. cit., the result 
is stated  under the extra hypothesis
that $r \ndiv N$ (apart from $r \neq 3$).
The referee to an earlier version of this article informed us
that our proposition (at least when $r>3$ and $f$ has integer Fourier coefficients)
follows from Proposition~5.5 in~\cite{yoo2} or
Theorem 3.1.3 in~\cite{ohta:eisenstein}, though the desired result
is not stated as such in a simple form in either to be quoted 
(one would have to deduce it 
after reading the notations, etc.). In any case, 
Theorem 3.1.3 of~\cite{ohta:eisenstein} uses 
Proposition 3.4.2 in loc. cit, in whose proof the author
says that ``this inductive step was inspired by the similar
argument of Proposition 3.5'' of the preprint version of~\cite{tor}; 
in the publised version~\cite{tor}, this became Proposition 3.6, which
we are suitably modifying below.
\end{rmk}

\begin{proof}  
By Lemma~\ref{lem:r}, if $r \divs N$, then $w_r = -1$, and we are done.
Thus we may assume henceforth that
$r \ndiv N$. 

The rest of the proof for $r > 3$ is a generalization of the proof of Proposition~3.6
of~\cite{tor}. We give the full details for two reasons: one is that
the case of $r = 3$ requires referring to the case $r >3$,
and the second is that 
the proof in loc. cit. had an error that we have fixed here
(see Remark~\ref{rmk3} below).

\comment{
If $r > 3$, then the result follows from the proof of 
Proposition~3.6 of~\cite{tor}  with the following changes:
replace $\Z/r\Z$ by~$\Of/\pp$, \mbox{$\bmod\ r$} by~\mbox{$\bmod \pp$}, 
Lemma~3.1 by our Lemma~\ref{lem:mazur0}, 
Lemma~3.2 by our Lemma~\ref{lem:dum}, and Lemma~3.5 by~\cite[Lemma~2.1.1]{ohta:eisenstein}.
}

Suppose, contrary to the conclusion of the Proposition, 
that for every  prime~$p$ 
that divides~$N$, we have $w_p = 1$. Then by Lemma~\ref{lem:dum},
for every  prime~$p$ 
that divides~$N$, we have $p \equiv -1 \bmod \pp$ 
(note that $p \neq r$ since $r \ndiv N$).

Following \cite[p.~77 and p.~70]{mazur:eisenstein},
by a holomorphic modular form in~$\omega^{\tensor 2}$
on~$\Gamma_0(N)$ defined over a ring~$R$,
we mean a modular form in the sense of
\cite[\S1.3]{katz:antwerp350} (see also \cite[\S~VII.3]{deligne-rapoport}).
Thus such an object is a rule which assigns to each pair
$(E_{/T},H)$, where $E$ is an elliptic curve over an $R$-scheme~$T$
and $H$ is a finite flat subgroup scheme of~$E_{/T}$ of order~$N$,
a section of~$\omega_{\scriptscriptstyle E_{/T}}^{\tensor 2}$,
where $\omega_{\scriptscriptstyle E_{/T}}$ is the sheaf of invariant differentials.
Since $r \ndiv N$, 
by~\cite[Lemma~1.3.5]{ohta:eisenstein}, 
if $g$ is a modular form of weight~$2$
on~$\Gamma_0(N)$ with coefficients in~$\Of/\pp$,
then there is a holomorphic modular form  in~$\omega^{\tensor 2}$
on~$\Gamma_0(N)$ defined over~$\Of/\pp$, which we will
denote $g \bmod \pp$,  such that the
$q$-expansion of~$g \bmod \pp$ agrees with the $q$-expansion
of~$g$ modulo~$\pp$. 

If $M$ is a postive integer, then let
us say that a holomorphic modular form~$g$ in~$\omega^{\tensor 2}$
on~$\Gamma_0(M)$ defined over~$\Of/\pp$ is {\em special at level~$M$}
if $a_n(g) \equiv \sigma(\frac{n}{(n,M)}) \prod_{p \mid M} 
(-1)^{{\rm ord}_p(n)} \bmod \pp$
for all positive integers~$n$.
Using Lemma~\ref{lem:mazur0} and the fact that $f$ is an eigenvector
for all the Hecke operators,
we see that $f \bmod \pp$ is special at level~$N$. 

Consider first the case where $r > 3$.

\noindent{\em Claim:} 
If $M$ is a square-free integer and 
$g$ is a holomorphic modular form in~$\omega^{\tensor 2}$
on~$\Gamma_0(M)$ 
defined over~$\Of/\pp$ that is special at level~$M$ and 
$s$ is a prime that divides~$M$, then there exists
a holomorphic modular form in~$\omega^{\tensor 2}$
on~$\Gamma_0(M/s)$ 
defined over~$\Of/\pp$ that is special at level~$M/s$
(which is also square-free).
\begin{proof}
By Proposition~\ref{prop:eis},
there is an Eisenstein series~$E$ which is an eigenvector
for all the Hecke operators, with $a_\ell(E) = \ell + 1$
for all primes $\ell \ndiv M$,
$a_p(E) = p$ 
for all primes~$p$ that divide~$M$ except $p = s$, and $a_s(E) = 1$; also
since $r > 3$, we can talk about $E \bmod \pp$.
Let $p_1, \ldots, p_t$ be the distinct primes that divide~$M/s$.
Then for any positive integer~$n$,
$$a_n(E) = \sigma \bigg(\frac{n}{(n,M)} \bigg) 
\prod_{i=1}^{t}{{p_i}^{{\rm ord}_{p_i}(n)}}\ .$$
Since
$p_i \equiv -1 \bmod \pp$ for $i = 1, \ldots, t$,
we see that $a_n( E) \equiv a_n( g) \bmod \pp$ if $n$ is coprime to~$s$, and thus 
$(E(q)-g(q)) \bmod \pp$ 
is a power series in $q^s$, i.e., 
there is an $h(q) \in (\Of/\pp)[[q]]$
with $ h(q^s)$ equal to $(E(q)-g(q)) \bmod \pp$.
By~\cite[Lemma~2.1.1]{ohta:eisenstein}, $h(q)$ is the $q$-expansion
of a 
holomorphic modular form, which we again denote~$h$, in~$\omega^{\tensor 2}$
on~$\Gamma_0(M/s)$ defined over~$\Of/\pp$.

Let $g' = h/2$.
We shall now show that $g'$ is special of level~$M/s$.
Let $n$ be a positive integer, $m' = \frac{n}{(n,s)}$,
and $e = {\rm ord}_s(n)$ (so $n = m' s^e$). 
Then considering that $m'$ is coprime to~$s$, we have
$a_n(E) = a_{m'}(E) a_{s^{e+1}}(E)$ since $E$ is an eigenfunction and 
$a_n(g) \equiv a_{m'}(g) a_{s^{e+1}}(g) \bmod \pp$ since $g$ is special. So
\begin{eqnarray}\label{eqn1}
a_n(h)&  = & a_n(E) - a_n(g) \nonumber\\
& \equiv &
a_{m'}(E) a_{s^{e+1}}(E) - a_{m'}(g) a_{s^{e+1}}(g) \bmod \pp \nonumber\\
& \equiv & a_{m'}(g) (a_{s^{e+1}}(E) - a_{s^{e+1}}(g))\bmod \pp,
\end{eqnarray}
where the last congruence follows since $a_{m'}(g) \equiv a_{m'}(E) \bmod \pp$,
considering that $m'$ is coprime to~$s$.

	From the definitions and the fact that \(s \mid \mid M\),  we find that
	\[
	a_{s^{e+1}}(E) = \sigma \bigg( {\frac{s^{e+1}}{s}} \bigg) \prod_{p \mid M, p\neq s} p^{\operatorname{ord}_{p}(s^{e+1})} = \sigma(s^{e})
	\]
	and
	\[
	a_{s^{e+1}}(g) \equiv \sigma \bigg( {\frac{s^{e+1}}{s}}\bigg) \prod_{p \mid M} (-1)^{\operatorname{ord}_{p}(s^{e+1})} \equiv  \sigma(s^{e})(-1)^{e+1} \bmod \pp. 
	\]

So we have
\begin{eqnarray}\label{eqn2}
a_{s^{e+1}}(E) - a_{s^{e+1}}(g) 
 \equiv 
 (1-(-1)^{e+1})\sigma(s^{e}) 
\bmod \pp .
\end{eqnarray}
We claim that 
$(1-(-1)^{e+1})\sigma(s^{e}) = 2 \sigma(s^{e})$: if 
$e \equiv 1 \bmod 2$, then $\sigma(s^e) = 0$, so the equality holds trivially,
and if $e \equiv 0 \bmod 2$, then $1-(-1)^{e+1} = 2$, so again the equality holds.
Putting this proven claim in~(\ref{eqn2}) and the result 
in~(\ref{eqn1}),
we get
\begin{eqnarray}\label{eqn4}
a_n(h) \equiv  a_{m'}(g) \cdot 2\sigma(s^e) 
\equiv 2 \sigma\bigg(\frac{m'}{(m',M)}\bigg)  \prod_{p \mid M} (-1)^{{\rm ord}_p(m')}  
\cdot \sigma(s^e) 
\bmod \pp, 
\end{eqnarray}
where the last congruence follows since $g$ is special at level~$M$.
Now since $n = m' s^e$, with $m'$ coprime to~$s$ and $s \ndiv (M/s)$, we have
\begin{eqnarray}\label{eqn5}
\sigma\bigg(\frac{m'}{(m',M)}\bigg)  \sigma(s^e)  
= \sigma\bigg(\frac{m' s^e}{(m',M)}\bigg) 
= \sigma\bigg(\frac{m' s^e}{(m' s^e,M/s)}\bigg) 
= \sigma\bigg(\frac{n}{(n,M/s)}\bigg) 
\end{eqnarray}
and 
\begin{eqnarray}\label{eqn6}
\prod_{p \mid M} (-1)^{{\rm ord}_p(m')} 
=  \prod_{p \mid M, \ p \neq s} (-1)^{{\rm ord}_p(m's^e)} 
= \prod_{p \mid (M/s)} (-1)^{{\rm ord}_p(n)}.
\end{eqnarray}
Using~(\ref{eqn5}) and~(\ref{eqn6}) in~(\ref{eqn4}), and
recalling that $g' = h/2$, we see that
\begin{eqnarray*}
a_n(g') \equiv 
\sigma\bigg(\frac{n}{(n,M/s)}\bigg) \prod_{p \mid (M/s)} (-1)^{{\rm ord}_p(n)}
\bmod \pp,
\end{eqnarray*}
i.e., $g'$ is special of level~$M/s$, as claimed.
\end{proof}

Starting with $f \bmod \pp$ (note that $r \ndiv N$), and repeatedly using the claim,
we see that there is 
a holomorphic modular form that is special of level~$1$, which is nontrivial
since the coefficient of~$q$ is~$1 \bmod \pp$ for a special form
(of any level). 
But by~\cite[Lemma II.5.6(a)]{mazur:eisenstein},
there are no nontrivial holomorphic modular forms
of level~$1$ in $\omega^{\tensor 2}$ defined over 
a field of characteristic other than~$2$ and~$3$.
This contradiction proves the case where $r > 3$.

\comment{
The strategy of the proof 
Proposition~3.6 of~\cite{tor} is that 
if the conclusion of that proposition is not true, then 
one can lower the level of $f \bmod \pp$ 
(the notation is as in loc. cit.)  using
a suitable Eisenstein series $E \bmod \pp$ obtained from 
Proposition~\ref{prop:eis} one prime at a time to
get a certain \mbox{$\bmod\ r$} form of level~$1$ in whose
$q$-expansion, the coefficient of~$q$ is~$1$,
which contradicts \cite[Lemma II.5.6(a)]{mazur:eisenstein}, which says that 
there are no nontrivial \mbox{$\bmod\ r$} forms
of level~$1$ defined over a field of characteristic other than~$2$ and~$3$.
}

Suppose now that $r=3$ and $\Of/\pp \isom \Z/3\Z$.
In the proof of the claim above, 
if $M$ is not a prime, then 
we have at least one prime $p \divs M$
where $a_p(E) = p$, and so by 
by part~(a) of Proposition~\ref{prop:eis}, we 
can talk of $E \bmod  \pp = E \bmod 3$ even though $r \not> 3$, and the 
conclusion of the claim holds. 
So if $N$ is not prime,
then starting with $f \bmod \pp$ (note that $r \ndiv N$), and repeatedly using the claim,
we see that there is 
a holomorphic modular form that is special of level a prime, call it~$p$. 
If $N$ were prime to start with, then we can take~$p=N$ to again
have a holomorphic modular form that is special of level the prime~$p$.

At this point, 
we follow a strategy that goes back to Mazur (see the proof of Prop.~II.14.1 on p.~114
in~\cite{mazur:eisenstein}), part of which is used in
the proof of Propostion 3.4.2 in~\cite{ohta:eisenstein}. For
the sake of being clear, rather than simply referring
to the sources above for the rest of the proof, we give the full details in our context.

Since $\Of/\pp \isom \Z/3\Z$ by hypothesis, 
$f \bmod \pp$ is a form defined over~$\Z/3\Z$. 
Suppose first that $p \equiv 1 \bmod 3$.
Then we can consider $E \bmod 3$ by part~(b) of Proposition~\ref{prop:eis}, 
and lower the level again to get a form of level~$1$ 
for whom the coefficient of~$q$ is~$1$.  The
constant term of this form is~$a_0(E)$ modulo~$3$, and
$a_0(E) = (p-1)/24$. If $p \not\equiv 1 \bmod 9$, then $a_0(E)$ is a unit \mbox{$\bmod\ 3$},
and by multiplying the form above by the inverse of this 
unit, we get
a \mbox{$\bmod\ 3$} form with constant term~$1$ and for which the coefficient of~$q$ is non-zero, which
contradicts \cite[Prop~II.5.6(c)(ii)]{mazur:eisenstein}.
If $p \equiv 1 \bmod 9$, then $a_0(E)$ is zero modulo~$3$, and this contradicts
\cite[Prop~II.4.10]{mazur:eisenstein}.

Finally, we are left with the case $p \equiv 2 \bmod 3$ (recall 
that we are assuming that $ r \ndiv N$, so $r \neq p$, 
and $r =3$, so $p\neq 3$). 
We consider the form $h= 3 E - \text{``}\times 3\text{''} f$ 
defined over~$\Z/9\Z$, with notation as in Lemma~3.3.2 in~\cite{ohta:eisenstein}
(where the construction $\text{``}\times 3\text{''}$ is attributed to~\cite[p.~86]{mazur:eisenstein}).
Since $a_0(f) = 0$,
we have $a_0(h) =  a_0(3E) = (1 - p)/8$ and since $p \equiv 2 \bmod 3$, this 
is a unit \mbox{$\bmod\ 9$}. 
Multiplying~$h$ by the inverse of this unit, we get 
a \mbox{$\bmod\ 9$} form  with $q$-expansion beginning with~$1$ 
and for which the coefficient of~$q$ is non-zero, which 
contradicts \cite[Lemma II.5.6(b)]{mazur:eisenstein} as $9  \ndiv 12$. 
\end{proof}

\begin{rmk}   \label{rmk3}
We could like to take the opportunity to point out a correction 
in the  proof of 
Proposition~3.6 of~\cite{tor}: in equation~(3.6) in loc. cit., the second line is 
not justified, and a priori need not be true. The proposition above
generalizes Proposition~3.6 of loc. cit., and so the reader 
should see the proof above instead of the one given in loc. cit.
\end{rmk}
\comment{
Using SAGE, we found that for any optimal elliptic curve~$E$ in 
Cremona's database~\cite{cremona:onlinetables} of elliptic curves (which includes all
curves with conductor less than~$130,000$), 
if $N$ is square-free and the odd part of~$E(\Q)_{\tor}$ is nontrivial, 
then $w_p = -1$ for at least one prime~$p$ that divides~$N$
(which indicates that the hypothesis that $r$ not divide~$3N$ in Proposition~\ref{prop:mazur2}
may be redundant).
The statement in the previous sentence is not true without the hypothesis
that $N$ is square-free. 
For example, 
the optimal elliptic curve~490a1 (note that $490= 2 \cdot 5 \cdot 7^2$) has a $3$-torsion point,
but $w_p = 1$ for all $p$ dividing~$490$. At the same time,
for all optimal elliptic curves in Cremona's database (of arbitrary conductor)
that have a $5$ or~$7$-torsion point, there is a prime~$p$ that divides the conductor
such that  $w_p=-1$. 

}

\begin{prop}\label{prop:red}
Suppose that $r \ndiv 2N$. If $A[\m]$ is reducible, then so is 
$\rho_{\m}$. 
\end{prop}
\begin{proof} We prove the contrapositive. 
So suppose $\rho_{\m}$ is irreducible. Then by~\cite[Prop.~5.2(b)]{ribet:modreps}, 
$J_0(N)[\m]$ is also irreducible. But $A[\m]$ is a subrepresentation 
of $J_0(N)[\m]$. So $A[\m]$ is either trivial or all of~$J_0(N)[\m]$; in either case, it
is irreducible, as was to be shown.
\end{proof}

\begin{proof}[Proof of Theorem~\ref{thm:main}]
Recall that we are assuming that $N$ is square-free, $r \ndiv 6N$, and 
either $\rho_{\m}$ or~$A[\m]$ is reducible. 
By Proposition~\ref{prop:red}, in any case, $\rho_{\m}$ is reducible. Thus all the results
in this section before Proposition~\ref{prop:red} apply.
We modify the proof of Theorem~1.2 of~\cite{tor} in Section~3 of loc. cit., by
replacing \mbox{$\bmod\ r$} by~\mbox{$\bmod\ \pp$}, Proposition~3.6 by our Proposition~\ref{prop:mazur2},
Lemma~3.2 by our Lemma~\ref{lem:dum}, 
Lemma~3.1 by our Lemma~\ref{lem:mazur0}, $\Z/r\Z$ by~$\Of/\pp$, Lemma~3.5  
by~\cite[Lemma~2.1.1]{ohta:eisenstein}, and
``$f$ is ordinary at~$r$'' by ``$f$ is ordinary at~$\pp$'', to conclude that~$A[\m] \cap C_E$
is nontrivial. The remaining conclusions of the theorem follow considering that
$C_E$ is rational, as $N$ is square-free. 
\end{proof}

\bibliographystyle{amsalpha}         

\bibliography{biblio}
\end{document}